The average math SAT score for students applying to Suffolk University last year was 510, with a standard deviation of 89. You do NOT need to check CLT here.

A. What score does an applicant need in order to be in the top 10% of math SAT scores?

B. What score does an applicant need in order to be at the 96th percentile of math SAT scores?

C. Between what score does an applicant need in order to be at the bottom 12% of math SAT sCores?
Answer #1

Hello Sir/ Mam

Given that:

Mean = 510.00
Standard Deviation = 89.00

(a)

Top 10% will have a corresponding z-score of 1.2816

Hence,

Required X = (89*1.2816) + 510 = 624.06

(b)

96th percentile will have a corresponding z-score of 1.751

Hence,

Required X = (89*1.751) + 510 = 665.81

(c)

12th percentile will have a corresponding z-score of -1.175

Hence,

Required X = (89*-1.175) + 510 = 405.43

Hence, the required score range is 0 to 405.43

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.